two cards are drawn without replacement from a standard deck of 52 playing cards. What is the probability of choosing a diamond and then, without replacement, heart? Answer need in both reduced fraction if possible and as a decimal number rounded to four decimal places.
Answer #1

Standard deck contains 13 diamond and 13 heart cards.

If first card is selected as diamond.

P(Diamond) = 13/52 = 1/4

Since first card is not replaced, there are 51 total cards left.

Probability of second card as heart is drawn from remaining 51 card = 13/51

So,

P(First diamond And then heart) = 1/4 * 13/51

= 13/204 ( OR in decimal 0.0637)